find the slopes and y-intercepts of the following
y=6
4x+9y=-1
y=3x+1\5

Answer by blwinbbbles(106)

the equation (one of them)for a line is:
y = mx + b with m being the slope and b being the y intercept..
if the line is horizontal ie: y = the same number for each pair
then the slope or m = 0
if the line is vertical ie: x = the same number for each pair..then the slope is undefined.
So let's solve the problems:
y = 6 this is a vertical line so the slope or m = zero and the y intercept (where the line crosses the y axis) is = 6

4x + 9y = -1...the easiest is to get this in linear form. So substract 4x from both sides
9y = -4x -1 now divide by 9
y = -4/9(x) - 1/9
so slope (m) = -4/9 and the y intercept (b) = -1/9
Last one:
y = 3x + 1/5 see if you can do this one yourself!!!!
hopes this helps